The reaction NO2(g) + CO(g) CO2(g) + NO(g) has a rate constant of 2.57 M−1∙s−1 at 701 K and 567 M−1∙s−1 at 895 K. Find the activation energy in kJ/mol
Answer #1

2-13A T T lo 8 (2.5フ 2.303 x f-1x0 2-343 = -Ea / 70/-895 o.019 2-343 = Fan 19y × 191 o. o 19 x 89 SX701 2. 343 Xo.olgXg95x7이